# https://dangeng.github.io/factorized_diffusion/

### v1

@@ 1 Introduction
We present a simple method for controlling the factors of such decompositions, allowing a user to generate images that are perceived differently under different viewing conditions, yet still globally coherent. We apply this approach to generating a variety of perceptual illusions (Fig. 1). (i) Inspired by the classic work of Oliva et al. [42] we generate hybrid images, whose interpretation changes with viewing distance, which we achieve by controlling the generated image’s low and high frequency components. A decomposition into three subbands allows us to produce hybrid images with three different prompts, which we refer to as triple hybrids. (ii) We generate color images whose appearance changes when they are viewed in grayscale, a phenomena that naturally occurs under dim lighting. We call these color hybrids, and achieve this by controlling image luminance separately from its color. (iii) Finally, we produce images that change appearance under motion blur, which is achieved by using a blur kernel to decompose an image. We refer to these as motion hybrids.
@@
We also show that we can take components from existing images, and generate the remaining components conditioned on text. This recovers a simple method to solve inverse problems, and is highly related to prior work on using diffusion models for solving inverse problems [54, 28, 8, 10, 9, 64, 34, 53]. We apply this technique to producing hybrid images from real images. Finally, we show that using certain decompositions with our method recovers prior techniques for spatial [2] and compositional [32] control over text prompts.
@@ 2 Related Work
Diffusion model control. Diffusion models are capable of both generating and editing images conditioned on text prompts. By modifying the reverse process [35, 18, 2, 68, 63], finetuning [67, 47], performing text inversion [52, 37, 26, 61, 65], swapping attention maps [23, 14, 59], supplying instructions [4], or using guidance [14, 44, 31, 1, 40, 21], modifying the style, location, and appearance of content in an image has become a relatively accessible task. Another line of work on compositional generation [12, 33, 2, 63] shows that diffusion models can generate images that conform to compositions of text prompts. Our work builds upon this, and shows that similar techniques can be applied to prompting individual components of an image to produce perceptual illusions. Our work is also similar to Wang et al. [63], in which a diffusion model is used to generate stacks of images that emulate a zooming video. However, we focus on generating only a single image that can be understood at multiple resolutions.
Computational optical illusions. Optical illusions are entertaining, but can also serve as windows into human and machine perception [24, 62, 19, 39, 27, 56, 20, 13]. As such, much work has gone into developing computational methods for generating optical illusions [16, 43, 22, 6, 42, 18, 58, 60, 5, 7]. In classic work, Oliva et al. introduced hybrid images [42], which are images that change their appearance depending on viewing distance or duration [49]. These images work by exploiting the multiscale processing of human perception [41, 50]. By aligning low frequency components of one image and high frequency components of another image, the observer perceives the image as the former when seen from far away and as the latter when seen up close. By contrast, our approach generates hybrid images from scratch with a diffusion model, as opposed to fusing together two existing images, and thus avoids manual alignment steps and the need to find appropriate images, and also leads to fewer artifacts.
@@ 3.4 Decompositions Considered
A final interesting decomposition is of the form ${\mathbf{x}}=\sum^{N}_{i}a_{i}{\mathbf{x}}$, for $\sum^{N}_{i}a_{i}=1$. Taking $a_{i}=\frac{1}{N}$ recovers the compositional diffusion method of Liu et al. [33], in which noise estimates are averaged to sample from conjunctions of multiple prompts.
@@ 3.5 Inverse Problems
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, and $\alpha_{t}$ is determined by the variance schedule. The argument of $f_{1}$ is a sample from the forward process, given the reference image—that is, a noisy version of ${\mathbf{x}}_{\text{ref}}$ with the correct amount of noise for timestep $t$. Essentially, we project ${\mathbf{x}}_{t}$ such that its first component matches that of ${\mathbf{x}}_{\text{ref}}$. This amounts to solving a (noiseless) inverse problem characterized by $\mathbf{y}=f_{1}({\mathbf{x}}$). Much work has gone into developing methods to solve inverse problems using diffusion models as priors, and this extension of our method can be viewed as a simplified version of prior work [28, 8, 64, 53, 34].
@@ 4.2 Other Decompositions
Figure 6: Motion Hybrids. We show additional motion hybrid results. These are images that change appearance when motion blurred. Here, the motion from upper left to bottom right.
@@ 4.3 Inverse Problems
As discussed in Sec. 3.5, we can modify our approach to solve inverse problems, resulting in a technique highly similar to prior work [64, 28, 10, 8, 9, 54, 34]. While previous work investigates using diffusion priors for solving problems such as colorization, inpainting, super-resolution, or phase retrieval, we apply the idea towards generating hybrid images from real images. Specifically, we take low or high frequency components from a real image and use our method to fill in the missing components, conditioned on a prompt. Results are shown in Fig. 1 and Fig. 8. We also provide colorization results in Appendix 0.J.
4.4 Limitations and Negative Impacts
@@
One major limitation of our method is that the success rate is relatively low. While our method can produce decent images consistently, very high quality images are rarer. This can be seen in Fig. 9 and Fig. 18, in which we visualize random samples for hybrid images, color hybrids, and motion hybrids. We attribute this fragility to the fact that our method produces images that are highly out-of-distribution for the diffusion model. In addition, there is no mechanism by which prompts associated with one component are discouraged from appearing in other components. Another failure case of our method is that the prompt for one component may dominate the generated image. Empirically, the success rate of our method can be improved by carefully choosing prompt pairs (see Appendix 0.I for additional discussion), or by manually tuning decomposition parameters, but we leave improving the robustness of our method in general to future work.
The ability to better control powerful image synthesis models opens numerous societal and ethical considerations. We apply our method to generating illusions, which in some sense seeks to deceive perception, possibly leading to applications in misinformation. We believe this and other concerns deserve further study and careful thought.
@@ 5 Conclusion
We thank Patrick Chao, Aleksander Holynski, Richard Zhang, Trenton Chang, Utkarsh Singhal, Huijie Zhang, Bowen Song, Jeongsoo Park, Jeong Joon Park, Jeffrey Fessler, Liyue Shen, Qing Qu, Antonio Torralba, and Alexei Efros for helpful discussions. Daniel is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 1841052.
@@ Appendix 0.F Hybrid Images with Latent Diffusion Models
We show hybrid images resulting from using our method with Stable Diffusion v1.5, a latent diffusion model, in Fig. 11. As can be seen the results are decent, but have significant artifacts, due to applying bandpass filters in the latent space. We find that pixel diffusion models produce much higher quality samples.
@@ Appendix 0.H Further Analysis of Factorized Diffusion
The first is now a linear combination of ${\mathbf{x}}_{t}$ and $\epsilon_{\theta}$, and the second adds in the noise $\mathbf{z}$. Our analysis then applies to just the update’ function.
@@ Appendix 0.I Choosing Prompts
Figure 13: Colorization. Our method can also be used to solve inverse problems, such as colorization. We show grayscale images that we wish to colorize on the left. The color component is then generated conditioned on the text prompts displayed.
@@ Appendix 0.J Colorization
We also show colorization results in Fig. 13, using our method as an inverse problem solver, as discussed in Sec. 3.5. Specifically, we use the color space decomposition introduced in Sec. 3.4. During diffusion model sampling we hold the grayscale component fixed to the grayscale component of a real image that we want to colorize, and generate the color component. Note that this is highly similar to prior work [54, 28, 9].

### v2

@@ 1 Introduction
We present a simple method for controlling the factors of such decompositions, allowing a user to generate images that are perceived differently under different viewing conditions, yet still globally coherent. We apply this approach to generating a variety of perceptual illusions (Fig. 1). (i) Inspired by the classic work of Oliva et al. [46] we generate hybrid images, whose interpretation changes with viewing distance, which we achieve by controlling the generated image’s low and high frequency components. A decomposition into three subbands allows us to produce hybrid images with three different prompts, which we refer to as triple hybrids. (ii) We generate color images whose appearance changes when they are viewed in grayscale, a phenomena that naturally occurs under dim lighting. We call these color hybrids, and make these by controlling image luminance separately from color. (iii) Finally, we produce images that change appearance under motion blur, by using a blur kernel to decompose an image. We refer to these as motion hybrids.
@@
We also show that we can take components from existing images, and generate the remaining components conditioned on text. This recovers a simple method to solve inverse problems, and is highly related to prior work on using diffusion models for solving inverse problems [58, 32, 11, 13, 12, 69, 38, 57, 9, 2]. We apply this technique to producing hybrid images from real images. Finally, we show that using certain decompositions with our method recovers prior techniques for spatial [4] and compositional [36] control over text prompts.
@@ 2 Related Work
Diffusion model control. Diffusion models are capable of both generating and editing images conditioned on text prompts. By modifying the reverse process [39, 21, 4, 73, 68], finetuning [72, 51], performing text inversion [56, 41, 30, 66, 70], swapping attention maps [26, 17, 64], supplying instructions [6], or using guidance [17, 48, 35, 3, 44, 24], modifying the style, location, and appearance of content in an image has become a relatively accessible task. Another line of work on compositional generation [15, 37, 4, 68] shows that diffusion models can generate images that conform to compositions of text prompts. Our work builds upon this, and shows that similar techniques can be applied to prompting individual components of an image to produce perceptual illusions. Our work is also similar to Wang et al. [68], in which a diffusion model is used to generate multiple images that blend seamlessly into a zooming video. However, we focus on generating only single images that can be understood at multiple resolutions.
Another line of work uses pretrained diffusion models to solve inverse problems, such as colorization or inpainting, in a zero-shot setting [58, 32, 11, 13, 12, 69, 38, 57, 9, 2]. We show that an extension of our method recovers these techniques, and we use it to generate hybrid images, which has not been considered before.
Computational optical illusions. Optical illusions are entertaining, but can also serve as windows into human and machine perception [27, 67, 22, 43, 31, 61, 23, 16]. As such, much work has gone into developing computational methods for generating optical illusions [19, 47, 25, 8, 46, 21, 63, 65, 7, 10]. In classic work, Oliva et al. introduced hybrid images [46], which are images that change their appearance depending on viewing distance or duration [53]. These images work by combining low frequencies from one image with high frequencies of another, exploiting the multiscale processing of human perception [45, 54]. By contrast, our approach generates hybrid images from scratch with a diffusion model. This avoids manual alignment steps and leads to higher quality illusions.
@@ 3.4 Decompositions Considered
A final interesting decomposition is of the form ${\mathbf{x}}=\sum^{N}_{i}a_{i}{\mathbf{x}}$, for $\sum^{N}_{i}a_{i}=1$. Taking $a_{i}=\frac{1}{N}$ recovers the compositional diffusion method of Liu et al. [37], in which noise estimates are averaged to sample from conjunctions of multiple prompts. Taking $a_{0}=1-\gamma$ and $a_{1}=\gamma$, gives us CFG [29] and negative prompting [1].
@@ 3.5 Inverse Problems
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, and $\alpha_{t}$ is determined by the variance schedule. The argument of $f_{1}$ is a sample from the forward process, given the reference image—that is, a noisy version of ${\mathbf{x}}_{\text{ref}}$ with the correct amount of noise for timestep $t$. Essentially, we project ${\mathbf{x}}_{t}$ such that its first component matches that of ${\mathbf{x}}_{\text{ref}}$. This amounts to solving a (noiseless) inverse problem characterized by $\mathbf{y}=f_{1}({\mathbf{x}}$). Much work has gone into developing methods to solve inverse problems using diffusion models as priors, and this extension of our method can be viewed as a simplified version of prior work [32, 11, 69, 57, 38, 9, 2].
@@ 4.2 Other Decompositions
Figure 6: Motion Hybrids. We show additional motion hybrid results. These are images that change appearance when motion blurred. Here, the motion is from upper left to bottom right.
@@ 4.3 Inverse Problems
As discussed in Sec. 3.5, we can modify our approach to solve inverse problems, resulting in a technique highly similar to prior work [69, 32, 13, 11, 12, 58, 38, 9, 2]. While previous work investigates using diffusion priors for solving problems such as colorization, inpainting, super-resolution, or phase retrieval, we apply the idea towards generating hybrid images from real images. Specifically, we take low or high frequency components from a real image and use our method to fill in the missing components, conditioned on a prompt. Results are shown in Fig. 1 and Fig. 8. We also provide colorization results in Appendix 0.J.
4.4 Limitations, Random Examples, and Societal Impacts
@@
While our method can produce decent images fairly consistently, very high quality images are rarer. This can be seen in Fig. 9 and Fig. 18, in which we visualize random samples for hybrid images, color hybrids, and motion hybrids. We attribute this to the fact that our method produces images that are highly out-of-distribution for the diffusion model. Another failure case of our method is that the prompt for one component may dominate the generated image. Empirically, the success rate of our method can be improved by carefully choosing prompt pairs (see Appendix 0.I for additional discussion), or by manually tuning decomposition parameters. Finally, our method relies of pixel diffusion models, as the image decompositions must be done in pixel space (see Appendices 0.A and 0.F for additional discussion and results).
The ability to better control image synthesis opens numerous societal and ethical considerations. We apply our method to generating illusions, which in a sense seeks to deceive perception, possibly leading to applications in misinformation. We believe this and other concerns deserve further study.
@@ 5 Conclusion
We thank Patrick Chao, Aleksander Holynski, Richard Zhang, Trenton Chang, Utkarsh Singhal, Huijie Zhang, Bowen Song, Jeongsoo Park, Jeong Joon Park, Jeffrey Fessler, Liyue Shen, Qing Qu, Antonio Torralba, and Alexei Efros for helpful discussions. We also thank Walter Scheirer, Luba Elliott, and Nicole Finn for reaching out and giving us the (amazing) opportunity to create an illusion for the CVPR 2024 T-shirt as part of the AI Art Gallery (see Appendix 0.L). Daniel is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 1841052.
@@ Appendix 0.F Hybrid Images with Latent Diffusion Models
We show hybrid images resulting from using our method with Stable Diffusion v1.5, a latent diffusion model, in Fig. 11. As can be seen the results are recognizable, but have significant artifacts, due to applying bandpass filters in the latent space. We find that pixel diffusion models produce much higher quality samples.
@@ Appendix 0.H Further Analysis of Factorized Diffusion
The first step is a linear combination of ${\mathbf{x}}_{t}$ and $\epsilon_{\theta}$, and the second adds in the noise $\mathbf{z}$. Our analysis then applies to just the update’ function.
@@ Appendix 0.I Choosing Prompts
Figure 13: Colorization. Our method can also be used to solve inverse problems, such as colorization. We show grayscale images that we wish to colorize on the left. The color component is then generated conditioned on the text prompts displayed. Note that this is effectively prior work [58, 32, 12].
@@ Appendix 0.J Colorization
We also show colorization results in Fig. 13, using our method as an inverse problem solver, as discussed in Sec. 3.5. Specifically, we use the color space decomposition introduced in Sec. 3.4. During diffusion model sampling we hold the grayscale component fixed to the grayscale component of a real image that we want to colorize, and generate the color component. Note that this is effectively prior work [58, 32, 12].
@@ Appendix 0.K Additional Results
Appendix 0.L CVPR 2024 T-Shirt Design
We created an inverse hybrid image for the official CVPR 2024 T-shirt as part of the AI Art track. Our goal was for attendees to see only a watercolor of the Seattle skyline when they received the shirt. Then, as they see other people wearing the shirts in the conference center from a distance, the text “CVPR” would be revealed.
We took an existing photo of the Seattle skyline, and pasted the text “CVPR” over the image. We then used our technique to condition an image on the low frequencies of the edited photo, and fill in the high frequencies given the text prompt ‘‘a watercolor of the seattle skyline with mount rainier in the background’’. The resulting image was then touched up by running Adobe Photoshop’s generative fill in a few locations with artifacts to improve quality. We show the low frequency image and the hybrid image, before editing, in Fig. 19. We also show additional candidate T-shirt designs in Fig. 20, which all reveal the text “CVPR” when viewed from a distance.
Figure 19: CVPR Hybrid Image T-shirt Design. a) The edited photo, from which we extract low spatial frequencies. b) The resulting hybrid image, after generating high spatial frequencies conditioned on the extracted low frequencies. For more details, please visit our website. Photo source: Pavol Svantner [60].
Figure 20: CVPR Hybrid Image T-shirt Design Candidates. We show more CVPR T-shirt designs. For easier viewing, we provide insets of each hybrid image at lower resolution. Best viewed digitally, with zoom.